OpenStudy (anonymous):

Why can't we apply L'Hospital's Rule to the lim as x->a f(x) ^ g(x) if f(x) = negative infinity and g(x) = 0 but it works if f(x) is infinity?

OpenStudy (anonymous):

for one thing if the base is negative you are in a lot of trouble the outputs will not be real numbers

OpenStudy (anonymous):

I don't believe L'Hospital's can be done here ... I think it only works when there is a quotient of two functions and either both tend to +/- infinity or both tend to zero.

OpenStudy (anonymous):

it is not a matter of l'hopital or no l'hopital you cannot have a negative base

OpenStudy (anonymous):

Unless natural logarithms are taken.

OpenStudy (anonymous):

not with real numbers

OpenStudy (anonymous):

what, for example, would \(-10^{\frac{1}{2}}\) be?

OpenStudy (anonymous):

the base of any exponential function has to be positive

OpenStudy (anonymous):

That would be imaginary. But they can still be raised to integer powers.

OpenStudy (anonymous):

logical reasoning, yes I get that part and taking the natural log satellite, well its not defined in real numbers for a negative base to certain fractional powers but it is defined for many other numbers like (-2)^2 or -10^(1/3)

OpenStudy (anonymous):

(-9)^(1/3) is -3 for example

OpenStudy (anonymous):

i am not saying you cannot take a negative number to some powers of your choosing i am saying that the base of any exponential function must be positive, or else it is not a real valued function there is no such function as for example \(f(x)=(-3)^x\)

OpenStudy (anonymous):

if you have a function with a variable in the exponent, that variable can take on infinitely many values, not just integers or specific fractions.

OpenStudy (anonymous):

in fact, how do we even interpret something like \(x^{\sin(x)}\)? the definition is given in terms of the logarithm, so \[x^{\sin(x)}=e^{\sin(x)\ln(x)}\] and is only defined for positive values of \(x\)

OpenStudy (anonymous):

thats a good point